# A Bayesian Generative Modeling Approach for Arbitrary Conditional Inference

### v1

@@ 2.5 Model Initialization
The parameters of neural net (e.g., weights and biases) are typically initialized by a uniform or normal distribution (Narkhede et al., 2022). Inspired from our previous works (Liu et al., 2024; Liu and Wong, 2025), the model performance can be further improved through an encoding generative modeling (EGM) initialization strategy. Specifically, an auxiliary pseudo-inverse encoder function $E$ is added to BGM to directly map the $X$ to the latent variable. Specifically, we desire that the distribution of $Z=E(V)$ should match the prior distribution of latent variable $Z$, which is a standard multivariate Gaussian distribution. The distribution match is achieved by adversarial training (Goodfellow et al., 2014). By the encoding process, the high-dimensional covariates with unknown distribution are mapped to a low-dimensional latent space with a desired distribution. Both the latent variable $Z$ and model parameters in $G$ are initialized through the EGM process.
@@ 2.6 Model Hyperparameters
For MNIST image data (Deng, 2012), The generative model $G$ adopts a fully convolutional decoder network that maps a 10-dimensional latent space to the $28\times 28$ image space. The latent vector $Z$ is first passed through a fully connected layer and reshaped into a low-resolution feature map of size $7\times 7\times 4F$, where $F$ is the number of base kernels (default $F=32$). The feature map is then progressively upsampled via two transposed convolution layers with stride 2, producing a $28\times 28$ spatial resolution, followed by an additional convolution layer for refinement. Each convolution block is equipped with batch normalization (Ioffe and Szegedy, 2015) and LeakyReLU activations. Two parallel $1\times 1$ convolution output heads generate the pixel-wise mean and variance, with Softplus applied to the variance head to ensure positivity.
We train BGM using the Adam optimizer (Kingma, 2014) with learning rate $0.005$ for both latent-variable updates and model-parameter updates. Training proceeds in mini-batches of size 32 for a default total of 500 epochs. To obtain a well-behaved initialization of both the latent space and model parameters, and to prevent early variance inflation, we adopt an EGM initialization as a warm start. This initialization phase uses up to 50,000 randomly sampled mini-batches prior to running the alternating stochastic iterative updating algorithm described in Section 2.3.
@@ 3.1 Convergence of the stochastic iterative updating algorithm
We introduce the main assumptions used in Stochastic approximation (Robbins and Monro, 1951) and nonconvex stochastic optimization (Ghadimi and Lan, 2013) as follows.
@@
The sequence $\{w_{t}\}$ remains almost surely in a compact set $\mathcal{W}\subset\mathbb{R}^{d_{z}+d_{\phi}}$. Denote full-data gradient of $\mathcal{J}$ by $\nabla\mathcal{J}=(\nabla_{Z}\mathcal{J},\nabla_{\phi}\mathcal{J})$ . $\mathcal{J}$ has $L$-Lipschitz gradient:
@@
The stochastic gradients satisfy unbiasedness: $\mathbb{E}\!\left[g_{t}^{(Z)}\mid\mathcal{F}_{t}\right]=\nabla_{Z}J(w_{t})$ and $\mathbb{E}\!\left[g_{t}^{(\phi)}\mid\mathcal{F}_{t}\right]=\nabla_{\phi}J(w_{t})$ where $\mathcal{F}_{t}$ is the filtration generated by the stochastic iterative updating algorithm up to iteration $t$. There exsits constant $\sigma_{Z}^{2}$ and $\sigma_{\phi}^{2}$ such that $\mathbb{E}\,\|g_{t}^{(Z)}|\mathcal{F}_{t}\|^{2}\leq\sigma_{Z}^{2}$ and $\mathbb{E}\,\|g_{t}^{(\phi)}|\mathcal{F}_{t}\|^{2}\leq\sigma_{\phi}^{2}$ for all $t$.
@@
Assumption 1 holds in the default architecture of generative model $G$ on compact sets where both Leaky-ReLU and Softplus activation functions are globally Lipschitz. Assumption 3 is a standard Robbins–Monro stepsize condition (Robbins and Monro, 1951) in stochastic approximation algorithms. Then we have the following theorems for convergence under above assumptions.
@@ 3.2 Statistical Consistency of the Learned Generative Model
Because neural network parameter space is unbounded and highly non-convex, uniform statistical control must be restricted to expanding but compact subsets of $\Phi$. Following standard practice in modern M-estimation (Shen and Wong, 1994; De Menezes et al., 2021), we work with a sieve sequence
@@ 3.3 Conditional‑Risk Bounds for Arbitrary Conditional Inference
where $k$ is the kernel function (e.g., RBF kernel), $y$ represents the observed value of the response variable and $r$ represents the predicted conditional distribution of $\mathcal{X_{B}}$ given an $x_{\mathcal{A}}$ (e.g., $P_{\mathcal{B}\mid\mathcal{A};\,\hat{\phi}_{N}}\!\left(\,\cdot\mid x_{\mathcal{A}}\right)$). The loss function is the squared Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) between the Dirac distribution at $y$ and the predictive distribution $r$ using kernel $k$.
@@ 4 Empirical Results
To demonstrate the performance of BGM algorithm, we conducted a series of empirical experiments based on both simulation datasets and real datasets. We benchmarked BGM against the state-of-the-art conformal prediction methods for conditional inference tasks. In the task of data imputation, we employed the MNIST, the handwritten digits image dataset to demonstrate that the imputation power offered by BGM. We also compared BGM to the widely used imputation methods to show the superior performance of BGM in imputation tasks.
@@ 4.2 Baseline Methods
For point estimation in conditional prediction tasks, we compare BGM method against Linear Regression, Random Forest (Breiman, 2001), XGboost (Chen, 2016), and a neural network predictor. The neural network architecture follows the default configuration recommended in localized conformal prediction (LCP) (Guan, 2023) as the default predictor.
For interval estimation in conditional prediction tasks, BGM is benchmarked against eight different conformal prediction (CP) methods for conditional inference. The conformal prediction provides finite-sample, distribution-free marginal coverage guarantees without imposing parametric assumptions on the data generation process. The performance of CP methods depends critically on the choice of nonconformity score computed on a calibration set. Our baselines include vanilla CP, which constructs prediction intervals using absolute regression residuals; locally weighted residual CP (LW-CP) (Lei et al., 2018), which normalizes residuals by an estimated local noise level to account for heteroscedasticity; quantile-regression–based CP (QR-CP)(Romano et al., 2019), which uses estimated lower and upper conditional quantiles to form tighter conformal scores, and locally weighted quantile CP (LWQR-CP), which further adjusts the quantile-based score using local variability estimates. We additionally incorporate recent advances in localized conformal prediction (LCP) (Guan, 2023), which extends CP to heterogeneous settings by weighting calibration samples according to their similarity to the test point. Applying localization to the above score constructions yields three additional baselines LW-LCP, QR-LCP, LWQR-LCP.
For data imputation tasks, we compared BGM to widely used imputation approaches, including mean imputation and MICE (Van Buuren and Groothuis-Oudshoorn, 2011). We provide implementation details of baseline methods in Appendix E.
@@ 4.3 Conditional Prediction
We generate $N=20,000$ observations as $\{(V^{i},R^{i})|i=1,...,N\}$. In the BGM setting, we have the fixed partition pattern as $X=(X_{\mathcal{A}},X_{\mathcal{B}})$ where $X_{\mathcal{A}}=V$, $X_{\mathcal{B}}=R$ and $p_{1}=d$, $p_{2}=1$. We randomly splited the data into 80% training set and 20% testing set. During training stage, we concatenated $V$ and $R$ from training set and fed the data to BGM for learning the joint distribution through the stochastic updating algorithm. In the testing stage, we inferred the posterior distribution of $P(R|V)$ on the held-out testing set and evaluated the model performance on both point estimate and interval estimate. For conformal prediction (CP) methods, 20% of the training set was further retained for calibration purposes. A neural network with three fully connected layers is used for prediction in all CP methods as suggested by (Guan, 2023). To evaluate the influence of the dimensionality of observation data $X$, we varied the dimension $p=d+1$ from 50, to 100, and 300. The significance level $\alpha$ is set to be 0.05.
Table 1: Comparison of point estimation performance across methods. Metrics include mean squared error (MSE), Pearson correlation (PCC), Spearman correlation (SCC) at different dimensions pp.
Dimension Metric Linear Regression Random Forest XGBoost LCP BGM
@@
In the point estimation experiments, BGM consistently demonstrates the best point prediction performance among all competing methods across three different evaluation metrics (Table 1). Traditional linear regression performs substantially worse in all settings, especially as the dimension increases. The neural network adopted from LCP is the best baseline, compared to other machine learning methods, across different settings. BGM further improves the best baseline by achieving a relative reduction in MSE ranging from 8.9% to 15.8% for different observation dimension. In addition, BGM achieves consistently higher correlation measures, improving PCC and SCC by 1.8% to 5.4% over the strongest baseline. These results highlight BGM’s superior ability to capture complex structures in high-dimensional conditional prediction tasks.
Table 2: Comparison of interval estimation performance across CP baselines and BGM. Metrics include Pearson correlation (PCC), Spearman correlation (SCC), empirical marginal coverage, and average prediction interval length (ave.PI) at different dimensions pp. The the nominal coverage level is set to $1-\alpha=0.95$ with $\alpha=0.05$. For reference, the average oracle interval lengths are 1.427 ($p=50$), 1.446 ($p=100$), and 1.400 ($p=300$).
@@
Next, we evaluate the interval estimation to quantify the ability of capturing the uncertainty in the prediction tasks. As shown in Table 2, BGM provides substantially better interval estimation than all conformal prediction (CP) baselines across all observational data dimensions. Vanilla CP fails to adapt to heteroscedastic noise because it relies on a single global residual quantile, resulting in nearly constant-width intervals across testing data points. Conformal prediction methods with localization largely improves the performance of standard CP by weighting calibration samples according a similarity-based localizer function relative to the test point, producing a local rather than global empirical distribution of nonconformity scores. Across all methods, BGM achieves the strongest alignment between the predicted and oracle interval lengths with the highest Pearson and Spearman correlations, ranging from 0.863 to 0.937 and 0.935 to 0.987, respectively, whereas the best CP baseline typically yields correlation between 0.6 and 0.9. In high dimensional setting (e.g., $p=300$), BGM substantially improves the PCC over the best CP baseline by 0.251, demonstrating its superior ability to adapt to underlying heteroscedasticity. In addition, BGM attains empirical coverage rate between 0.944 and 0.966, close to the nominal 95% level, while all CP baselines are systematically more conservative with coverage from 0.980 to 0.991 but with considerably wider prediction intervals.
@@
To further examine the quality of the estimated prediction intervals, we compare BGM with the top three CP baselines by plotting the predicted interval lengths against the oracle interval lengths for all test points (Figure 2). Here, oracle prediction interval is obtained via Monte Carlo integration under the true data-generating process. Specifically, for each test point ($v^{i}$), we first sample $Z^{i}\sim p(Z\mid V=v^{i})$ using its closed-form Guassian posterior. We draw $M$ Monte Carlo samples $\{Z^{i,m}\mid m=1,..,M\}$ of the latent variables, followed by sampling $R^{i,m}\sim p(R\mid Z=z^{i,m})$. The oracle interval is defined by the empirical $\alpha/2$- and (1-$\alpha/2$) quantiles of $\{R^{i,m}\mid m=1,..,M\}$. These scatter plots provide a visual assessment of calibration across different methods. Across all settings, BGM exhibits the strongest alignment with the oracle intervals, yielding substantially higher Pearson correlations than the competing methods, particularly in higher-dimensional settings.
Overall, these results demonstrate that BGM not only provides accurate point estimates in the conditional prediction tasks but also achieves more accurate interval estimates than existing CP approaches with closer-to-oracle calibration of uncertainty. Together, this highlights BGM’s ability to deliver reliable and efficient predictions with uncertainty qualifications, even in challenging high-dimensional and heteroscedastic settings.
@@ 4.4 Data Imputation
We further investigate the ability of BGM to impute missing values using the MNIST handwritten digits dataset. Each image is represented as a $28\times 28$ grayscale intensity vector, rescaled to $[0,1]$ (Figure 3a). We train an unconditional BGM model on the full MNIST training set with $60,000$ images, where the generative function $G$ is represented by a convolutional decoder network. After training, BGM learns the joint distribution of all pixels. We generated MNIST images with BGM by first randomly sampling from the prior distribution of the latent space and then sampling the mean and variance through the learned $G$ function. The generated MNIST images well assemble the distribution of the true images (Figure 3b).
@@
To quantitatively assess the utility of BGM imputations for downstream prediction, we trained a standard convolutional neural network (CNN) classifier based on the original, fully observed MNIST training set. On clean test images this classifier achieves an accuracy of $0.9914$. We then evaluate the classification performance under six levels of increasing missingness on the testing set. Classification accuracy deteriorates rapidly as more pixels are removed, dropping from $0.9513$ to $0.6560$ as missingness increases. We next compared several imputation strategies applied prior to classification. As shown in Figure 4a, BGM imputations consistently provide the most substantial accuracy improvement across all masking levels, achieving accuracies of 0.966% to 0.988% and outperforming the classical imputation baselines.
@@
Figure 4: Data imputation with uncertainty qualification on MNIST dataset. (a) Imputation improves MNIST test classification accuracy by different methods. We varied the level of missingness (e.g., number of random $5\times 5$ holes). (b-f) BGM imputed images and uncertainty heatmap with different missing patterns. (b) A $5\times 13$ stripe mask in the middle. (c) A $5\times 5$ square mask in the upper left. (d) A $5\times 5$ square mask in the upper right (e) A $5\times 5$ square mask in the lower left. (f) A $5\times 5$ square mask in the lower right. Note that the pixels within the green rectangles are imputed by BGM posterior mean and the uncertainty is calculated by the average prediction interval length with $\alpha=0.05$ across all testing images.
These results demonstrate that BGM learns a realistic conditional distribution over observed image pixels and can act as a powerful plug-in data imputer. Moreover, BGM offers the full posterior distribution for the missing data imputations, providing much richer information compared to traditional data imputation methods and more flexibility for the missing patterns.

### v2

@@ 2.5 Model Initialization
The parameters of neural net (e.g., weights and biases) are typically initialized by a uniform or normal distribution (Narkhede et al. 2022). Inspired from our previous works (Liu et al. 2024, Liu and Wong 2025), the model performance can be further improved through an encoding generative modeling (EGM) initialization strategy. Specifically, an auxiliary pseudo-inverse encoder function $E$ is added to BGM to directly map the $X$ to the latent variable. Specifically, we desire that the distribution of $Z=E(V)$ should match the prior distribution of latent variable $Z$, which is a standard multivariate Gaussian distribution. The distribution match is achieved by adversarial training (Goodfellow et al. 2014, Liu et al. 2021). By the encoding process, the high-dimensional covariates with unknown distribution are mapped to a low-dimensional latent space with a desired distribution. Both the latent variable $Z$ and model parameters in $G$ are initialized through the EGM process. After initialization, the encoder function $E$ is removed during model training process.
@@ 2.6 Model Hyperparameters
For MNIST image data (Deng 2012), The generative model $G$ adopts a fully convolutional decoder network that maps a 10-dimensional latent space to the $28\times 28$ image space. The latent vector $Z$ is first passed through a fully connected layer and reshaped into a low-resolution feature map of size $7\times 7\times 4F$, where $F$ is the number of base kernels (default $F=32$). The feature map is then progressively upsampled via two transposed convolution layers with stride 2, producing a $28\times 28$ spatial resolution, followed by an additional convolution layer for refinement. Each convolution block is equipped with batch normalization (Ioffe and Szegedy 2015) and LeakyReLU activations. Two parallel $1\times 1$ convolution output heads generate the pixel-wise mean and variance, with Softplus applied to the variance head to ensure positivity.
We train BGM using the Adam optimizer (Kingma 2014) with learning rate $0.005$ for both latent-variable updates and model-parameter updates. Training proceeds in mini-batches of size 32 for a default total of 500 epochs. To obtain a well-behaved initialization of both the latent space and model parameters, and to prevent early variance inflation, we adopt an EGM initialization as a warm start. This initialization phase uses up to 50,000 randomly sampled mini-batches prior to running the alternating stochastic iterative updating algorithm described in Section 2.3.
@@ 3.1 Convergence of the stochastic iterative updating algorithm
We introduce the main assumptions used in Stochastic approximation (Robbins and Monro 1951) and nonconvex stochastic optimization (Ghadimi and Lan 2013) as follows.
@@
The sequence $\{w_{t}\}$ remains almost surely in a compact set $\mathcal{W}\subset\mathbb{R}^{d_{z}+d_{\phi}}$. Denote full-data gradient of $\mathcal{J}$ by $\nabla\mathcal{J}=(\nabla_{Z}\mathcal{J},\nabla_{\phi}\mathcal{J})$ . $\mathcal{J}<\infty$ and $\mathcal{J}$ has $L$-Lipschitz gradient:
@@
The stochastic gradients satisfy unbiasedness: $\mathbb{E}\!\left[g_{t}^{(Z)}\mid\mathcal{F}_{t}\right]=\nabla_{Z}J(w_{t})$ and $\mathbb{E}\!\left[g_{t}^{(\phi)}\mid\mathcal{F}_{t}\right]=\nabla_{\phi}J(w_{t})$ where $\mathcal{F}_{t}$ is the filtration generated by the stochastic iterative updating algorithm up to iteration $t$. There exists constant $\sigma_{Z}^{2}$ and $\sigma_{\phi}^{2}$ such that $\mathbb{E}\,\|g_{t}^{(Z)}|\mathcal{F}_{t}\|^{2}\leq\sigma_{Z}^{2}$ and $\mathbb{E}\,\|g_{t}^{(\phi)}|\mathcal{F}_{t}\|^{2}\leq\sigma_{\phi}^{2}$ for all $t$.
@@
Assumption 1 holds in the default architecture of generative model $G$ on compact sets where both Leaky-ReLU and Softplus activation functions are globally Lipschitz. Assumption 3 is a standard Robbins–Monro stepsize condition (Robbins and Monro 1951) in stochastic approximation algorithms, which can be satisfied by an optimizer with designed decaying learning rate. Then we have the following theorems for convergence under above assumptions.
@@ 3.2 Statistical Consistency of the Learned Generative Model
Because neural network parameter space is unbounded and highly non-convex, uniform statistical control must be restricted to expanding but compact subsets of $\Phi$. Following standard practice in modern M-estimation (Shen and Wong 1994, De Menezes et al. 2021), we work with a sieve sequence
@@ 3.3 Conditional‑Risk Bounds for Arbitrary Conditional Inference
where $k$ is the kernel function (e.g., RBF kernel), $y$ represents the observed value of the response variable and $r$ represents the predicted conditional distribution of $\mathcal{X_{B}}$ given an $x_{\mathcal{A}}$ (e.g., $P_{\mathcal{B}\mid\mathcal{A};\,\hat{\phi}_{N}}\!\left(\,\cdot\mid x_{\mathcal{A}}\right)$). The loss function is the squared Maximum Mean Discrepancy (MMD) (Gretton et al. 2012) between the Dirac distribution at $y$ and the predictive distribution $r$ using kernel $k$.
@@ 4 Empirical Results
To demonstrate the performance of BGM algorithm, we conducted a series of empirical experiments based on both simulation and real datasets. We benchmarked BGM against the leading conformal prediction methods for conditional inference tasks. In the task of data imputation, we employed the MNIST, the handwritten digits image dataset to demonstrate that the imputation power offered by BGM. We also compared BGM to the widely used imputation methods to show the superior performance of BGM in imputation tasks.
@@ 4.2 Baseline Methods
For point estimation in conditional prediction tasks, we compare BGM method against Linear Regression, Random Forest (Breiman 2001), XGboost (Chen 2016), VAEAC (Ivanov et al. 2019) and a neural network predictor. The neural network architecture follows the default configuration recommended in localized conformal prediction (LCP) (Guan 2023) as the default predictor. Similarly to BGM, we used the conditional mean of multiple imputations from VAEAC as the point estimate where the number of imputations equals to the number of posterior sample size of BGM (default: 5000).
For interval estimation in conditional prediction tasks, BGM is benchmarked against eight different conformal prediction (CP) methods (Lei et al. 2018, Romano et al. 2019, Guan 2023) and generative AI method VAEAC for conditional inference. The conformal prediction provides finite-sample, distribution-free marginal coverage guarantees without imposing parametric assumptions on the data generation process. The performance of CP methods depends critically on the choice of nonconformity score computed on a calibration set. Similar to BGM, we constructed prediction intervals of VAEAC by taking the $\alpha$/2- and (1-$\alpha$/2)-quantiles based on multiple imputed predictions.
For data imputation tasks, we compared BGM to widely used imputation approaches, including mean imputation and MICE (Van Buuren and Groothuis-Oudshoorn 2011). We provide implementation details of baseline methods in Appendix E.
@@ 4.3 Conditional Prediction
We generate $N=20,000$ observations as $\{(V^{i},R^{i})|i=1,...,N\}$. In the BGM setting, we have the partition pattern as $X=(X_{\mathcal{A}},X_{\mathcal{B}})$ where $X_{\mathcal{A}}=V$, $X_{\mathcal{B}}=R$ and $p_{1}=d$, $p_{2}=1$. We randomly split the data into 80% training set and 20% testing set. During training stage, we concatenated $V$ and $R$ from training set and fed the data to BGM for learning the joint distribution through the stochastic updating algorithm. In the testing stage, we inferred the posterior distribution of $P(R|V)$ on the held-out testing set and evaluated the model performance on both point estimate and interval estimate. For conformal prediction (CP) methods, 20% of the training set was further retained for calibration purposes. A neural network with three fully connected layers is used for prediction in all CP methods as suggested by (Guan 2023). To evaluate the influence of the dimensionality of observation data $X$, we varied the dimension $p=d+1$ from 50, to 100, and 300. The significance level $\alpha$ is set to be 0.05.
Table 1: Comparison of point estimation performance across different methods. Metrics include mean squared error (MSE), Pearson correlation (PCC), Spearman correlation (SCC) at different dimensions pp.
Dimension Metric Linear Regression Random Forest XGBoost VAEAC LCP BGM
@@
In the point estimation experiments, BGM consistently demonstrates the best point prediction performance among all competing methods across three different evaluation metrics (Table 1). Traditional linear regression performs substantially worse in all settings, especially as the dimension increases. VAEAC performs well when $p=50$ and performance decreases significantly when $p=300$. The neural network adopted from LCP is the most robust baseline compared to other methods. BGM further improves the best baseline by achieving a relative reduction in MSE ranging from 6.2% to 14.6% for different observation dimension. In addition, BGM achieves consistently higher correlation measures (both PCC and SCC) over the strongest baseline. These results highlight BGM’s superior ability to capture complex structures in high-dimensional conditional prediction tasks.
Table 2: Comparison of interval estimation performance across BGM and baseline methods. Metrics include Pearson correlation (PCC), Spearman correlation (SCC), empirical marginal coverage, and average prediction interval length (ave.PI) at different dimensions pp. The the nominal coverage level is set to $1-\alpha=0.95$ with $\alpha=0.05$. For reference, the average oracle interval lengths are 1.427 ($p=50$), 1.446 ($p=100$), and 1.400 ($p=300$).
@@
Next, we evaluate the interval estimation to quantify the ability of capturing the uncertainty in the prediction tasks. As shown in Table 2, BGM provides substantially better interval estimation than all conformal prediction (CP) baselines across all observational data dimensions. Vanilla CP fails to adapt to heteroscedastic noise because it relies on a single global residual quantile, resulting in nearly constant-width intervals across testing data points. Conformal prediction methods with localization largely improves the performance of standard CP by weighting calibration samples according a similarity-based localizer function relative to the test point, producing a local rather than global empirical distribution of nonconformity scores. Across all methods, BGM achieves the strongest alignment between the predicted and oracle interval lengths with the highest Pearson and Spearman correlations, ranging from 0.863 to 0.937 and 0.935 to 0.987, respectively, whereas the best CP baseline typically yields correlation between 0.6 and 0.9. In high dimensional setting (e.g., $p=300$), BGM substantially improves the PCC over the best CP baseline by 0.251, demonstrating its superior ability to adapt to underlying heteroscedasticity while distance-based localization of CP methods fails. In addition, BGM attains empirical coverage rate between 0.944 and 0.966, close to the nominal 95% level, while all CP baselines are systematically more conservative with coverage from 0.980 to 0.991 but with considerably wider prediction intervals. We also benchmarked the performance of VAEAC in this task to demonstrate its inability to capture uncertainty (Table S1).
To further examine the quality of the estimated prediction intervals, we compare BGM with the top three CP baselines by plotting the predicted interval lengths against the oracle interval lengths for all test points (Figure 2). Here, oracle prediction interval is obtained via Monte Carlo integration under the true data-generating process. Specifically, for each test point ($v^{i}$), we first sample $Z^{i}\sim p(Z\mid V=v^{i})$ using its closed-form Gaussian posterior. We draw $M$ Monte Carlo samples $\{Z^{i,m}\mid m=1,..,M\}$ of the latent variables, followed by sampling $R^{i,m}\sim p(R\mid Z=z^{i,m})$. The oracle interval is defined by the empirical $\alpha/2$- and (1-$\alpha/2$) quantiles of $\{R^{i,m}\mid m=1,..,M\}$. These scatter plots provide a visual assessment of calibration across different methods. Across all settings, BGM exhibits the strongest alignment with the oracle intervals, yielding substantially higher Pearson correlations than the competing methods, particularly in higher-dimensional settings.
Overall, these results demonstrate that BGM not only provides accurate point estimates in the conditional prediction tasks but also achieves more accurate interval estimates than existing CP approaches with closer-to-oracle calibration of uncertainty. Together, this highlights BGM’s ability to deliver reliable and efficient predictions with uncertainty quantification, even in challenging high-dimensional and heteroscedastic settings.
@@ 4.4 Data Imputation
We further investigate the ability of BGM to impute missing values using the MNIST handwritten digits dataset. Each image is represented as a $28\times 28$ grayscale intensity vector, rescaled to $[0,1]$ (Figure 3a). We train an BGM model on the MNIST training set with $60,000$ images, where the generative function $G$ is represented by a convolutional decoder network. After training, BGM learns the joint distribution of all pixels. We generated MNIST images with BGM by first randomly sampling from the prior distribution of the latent space, which is a standard Gaussian distribution. Then we obtained the mean and variance through the learned $G$ function. The generated MNIST images using the learned mean and variance functions well assemble the distribution of the true images (Figure 3b).
@@
Figure 4: Data imputation on MNIST dataset. (a) Imputation improves MNIST test classification accuracy by different methods with different level of missingness (e.g., number of random $5\times 5$ holes). (b-f) BGM imputed images and uncertainty heatmap with different missing patterns. (b) A $5\times 13$ stripe mask in the middle. (c) A $5\times 5$ square mask in the upper left. (d) A $5\times 5$ square mask in the upper right (e) A $5\times 5$ square mask in the lower left. (f) A $5\times 5$ square mask in the lower right. Note that the pixels within the green rectangles are imputed by BGM posterior mean and the uncertainty is denoted by the average prediction interval length with $\alpha=0.05$ across all test images.
To quantitatively assess the utility of BGM imputations for downstream prediction, we trained a standard convolutional neural network (CNN) classifier based on the original, fully observed MNIST training set. On clean test images this classifier achieves an accuracy of $0.9914$. We then evaluate the classification performance under six levels of increasing missingness on the testing set. Classification accuracy deteriorates rapidly as more pixels are removed, dropping from $0.9513$ to $0.6560$ as missingness increases. We next compared several imputation strategies applied prior to classification. As shown in Figure 4a, BGM imputations consistently provide the most substantial accuracy improvement across all masking levels, achieving accuracies of $0.966$ to $0.988$ and outperforming the classical imputation baselines.
@@
These results demonstrate that BGM learns a realistic conditional distribution over observed image pixels and can act as a powerful plug-in data imputer. Moreover, BGM offers the full posterior distribution for missing data imputations, providing much richer information and more flexibility for the missing patterns compared to traditional data imputers.